what is the average of the other three numbers, if the average of six numbers is 12. three of the numbers are 8,9 and 16. please answer

Answer by fcabanski(1391)

Average = sum/(# of numbers)


For the six numbers, the problem gives (8+9+16 + (sum of other three))/6 = 12


8+9+16= 33 and call the Sum of the other three numbers S.


33 + S = 72


S= 72-33=39


Apply the average formula. S is the sum of the three numbers, and there are 3. S/3 = average. 39/3 = 13 is the average of the other three numbers.